# On three recent questions of Bourin and Lee on quadratic symmetric modulus and Euler-operator identity

### v1

@@ Abstract.
We answer three questions posed by Bourin and Lee on symmetric moduli and related orbit inequalities in [6]. First, we show that the exponent 22 in their unitary-orbit estimate for the quadratic symmetric modulus is optimal in every dimension $n\geq 2$ by an explicit $2\times 2$ counterexample for $p>2$. Second, we construct a compact operator for which the corresponding singular-value inequality fails for every $p>2$ (indeed for all $p>0$ at a fixed pair of indices). Finally, we obtain isometry-orbit refinements of Euler’s quadrilateral identity for matrices and derive several Clarkson–McCarthy type inequalities for Schatten $p$-norms and related consequences.
@@
Symmetric modulus; singular values; unitary orbits; Euler identity; Schatten norms
@@ 1. Introduction
Let $\mathcal{H}$ be a complex Hilbert space, and let $\mathbb{K}$ denote the ideal of compact operators on $\mathcal{H}$. For $Z\in\mathbb{K}$, the singular values $\{\mu_{m}^{\downarrow}(Z)\}_{m\geq 1}$ are defined as the eigenvalues (counted with multiplicity) of $|Z|=(Z^{*}Z)^{1/2}$, listed in nonincreasing order. Let $\mathbb{M}_{m,n}$ be the set of all complex $m\times n$ matrices; when $m=n$, we write $\mathbb{M}_{n}$. For a matrix $X\in\mathbb{M}_{n}$, we adopt the compatible convention that $\mu_{m}^{\downarrow}(X)=0$ for all $m>n$, so that $\{\mu_{m}^{\downarrow}(X)\}_{m\geq 1}$ may be viewed as an infinite nonincreasing sequence (and this agrees with the compact-operator notation).
We write
$\operatorname{Re}Z=\frac{Z+Z^{*}}{2},\qquad\operatorname{Im}Z=\frac{Z-Z^{*}}{2i}.$
For two Hermitian matrices $A,B\in\mathbb{M}_{n}$, we write $A\leq B$ if BAB-A is positive semidefinite.
For $p>0$, the Schatten pp-(quasi)norm on $\mathbb{M}_{n}$ is
$\|X\|_{p}:=\big(\operatorname{Tr}|X|^{p}\big)^{1/p},\qquad X\in\mathbb{M}_{n}.$
When $1<p<\infty$, we denote by $q$ the conjugate exponent, i.e. $1/p+1/q=1$.
Bourin and Lee [6, Corollary 4.3] obtained the following unitary-orbit estimate for the quadratic symmetric modulus.
@@
They also posed the following question concerning the optimality of the exponent 22.
Question 1.2 ([6, Question 4.5]).
@@
Our first result answers Question 1.2 affirmatively in every noncommutative dimension $n\geq 2$ by constructing a $2\times 2$ counterexample for $p>2$.
@@
For $n=1$, the statement of Question 1.2 is different: for a complex number $z$, the left-hand side of (1.1) equals $|z|$ for every $p>0$, while the right-hand side equals $|\Re z|+|\Im z|\geq|z|$. Hence (1.1) holds for all $p>0$ in dimension 11. The necessity $p\leq 2$ is therefore a genuinely noncommutative phenomenon requiring $n\geq 2$.
@@
Question 1.5 ([6, Question 4.7]).
@@
$\mu_{1+j+k}^{\downarrow}\!\left(\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\right)>\mu_{1+j+k}^{\downarrow}(\operatorname{Re}Z)+\mu_{1+k}^{\downarrow}(\operatorname{Im}Z)$
@@
We answer this question in the affirmative; in fact, we produce an example where the same pair of indices works for all $p>0$.
@@
There exists $Z\in\mathbb{K}$ such that for every $p>2$ one can find integers $j,k\geq 0$ for which
$\mu^{\downarrow}_{1+j+k}\!\left(\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\right)>\mu^{\downarrow}_{1+j+k}(\operatorname{Re}Z)+\mu^{\downarrow}_{1+k}(\operatorname{Im}Z).$ (1.2)
Moreover, one may take $j=k=1$, in which case the inequality holds for all $p>0$.
@@
We recall that an isometry is a rectangular matrix $W\in\mathbb{M}_{m,n}$ such that $W^{*}W=I_{n}$. Bourin and Lee [6] suggested exploring isometry/unitary-orbit refinements of Euler-type identities. Our following main results give isometry-orbit refinements of (1.3).
@@
Based on results obtained by the author in [8], we obtain the following two Clarkson–McCarthy type inequalities.
@@
$\|A+B\|_{p}^{p}+\|B+C\|_{p}^{p}+\|C+A\|_{p}^{p}$ $\leq 2^{p-2}\big(\|A+B+C\|_{p}^{p}+\|A\|_{p}^{p}+\|B\|_{p}^{p}+\|C\|_{p}^{p}\big),$
$\|A+B+C\|_{p}^{p}+\|A\|_{p}^{p}+\|B\|_{p}^{p}+\|C\|_{p}^{p}$ $\leq 3^{\frac{p}{2}-1}\big(\|A+B\|_{p}^{p}+\|B+C\|_{p}^{p}+\|C+A\|_{p}^{p}\big).$
For $0<p\leq 2$, both inequalities reverse.
Corollary 1.13.
Let $A,B,C\in\mathbb{M}_{n}$. Then for $1<p\leq 2$ ,
where $q$ is the conjugate exponent of pp. For $p\geq 2$, the inequality is reversed.
Remark 1.14.
Set C=BC=-B in Corollary 1.12 and Corollary 1.13, we derive the classic Clarkson–McCarthy inequality in [7].
Organization of this paper. In Section 2, we prove Theorem 1.3. In Section 3, we prove Theorem 1.6. In Section 4, we prove Theorems 1.9, 1.10, 1.11 and derive further related consequences.

### v2

# On five questions of Bourin and Lee: symmetric moduli and an Euler operator identity

@@ Abstract.
We answer five questions posed by Bourin and Lee on symmetric moduli and related orbit inequalities in [12], and thereby obtain a sequence of sharp results for matrices and compact operators. We first show that the isometry-orbit identity behind the matrix weighted parallelogram law cannot be extended beyond the parameter range $0\leq x\leq 1$, and that a counterexample already exists in dimension one. Next, we prove that the exponent $2$ in the Bourin–Lee unitary-orbit estimate for the quadratic symmetric modulus is optimal in every dimension $n\geq 2$ by constructing an explicit $2\times 2$ counterexample for all $p>2$. We then construct a compact operator $Z$ for which the associated singular-value inequality fails for every $p>2$, in fact for a fixed choice of indices. We also settle a Thompson-type triangle problem for symmetric moduli: the inequality fails for the arithmetic symmetric modulus but holds for the quadratic symmetric modulus. Finally, we develop isometry-orbit refinements of an Euler operator identity and derive sharp Clarkson–McCarthy type inequalities for Schatten $p$-norms, together with further consequences for unitarily invariant norms and singular values.
@@
Symmetric modulus; unitary orbits; Euler identity; Thompson inequality; Clarkson-McCarthy type inequalities
@@ 1. Introduction
Throughout this paper, $\mathcal{H}$ denotes a complex Hilbert space, and $\mathbb{K}$ the ideal of compact operators on $\mathcal{H}$. For $Z\in\mathbb{K}$, we write $\{\mu_{m}^{\downarrow}(Z)\}_{m\geq 1}$ for the singular values of $Z$, namely the eigenvalues of $|Z|=(Z^{*}Z)^{1/2}$ (counted with multiplicity) arranged in nonincreasing order. Let $\mathbb{M}_{m,n}$ be the space of complex $m\times n$ matrices, and write $\mathbb{M}_{n}:=\mathbb{M}_{n,n}$. Let $X\in\mathbb{M}_{n}$. We adopt the convention that $\mu_{m}^{\downarrow}(X)=0$ for all $m>n$, so that $\{\mu_{m}^{\downarrow}(X)\}_{m\geq 1}$ can be regarded as an infinite nonincreasing sequence, consistent with the compact-operator notation. We denote by $X^{\mathsf{T}}$ its transpose, and by $X^{*}$ its conjugate transpose (adjoint). For scalars $d_{1},\dots,d_{r}\in\mathbb{C}$, we denote by $\operatorname{diag}(d_{1},\dots,d_{r})$ the $r\times r$ diagonal matrix whose diagonal entries are $d_{1},\dots,d_{r}$ (and all off-diagonal entries are 0). More generally, if $X_{1}\in\mathbb{M}_{n_{1}},\dots,X_{k}\in\mathbb{M}_{n_{k}}$, then $\operatorname{diag}(X_{1},\dots,X_{k}):=X_{1}\oplus\cdots\oplus X_{k}$ denotes the block-diagonal matrix with diagonal blocks $X_{1},\dots,X_{k}$.
1.1. A problem on matrix weighted parallelogram law
For $A,B\in\mathbb{M}_{n}$ and $x\in\mathbb{R}$, we write $A\nabla_{x}B:=(1-x)A+xB$ for the weighted arithmetic mean of $A$ and $B$. We begin by recalling a unitary-orbit refinement of the weighted parallelogram identity due to Bourin and Lee [12].
Theorem 1.1 ([12, Corollary 2.2]).
Let $A,B\in\mathbb{M}_{n}$ and $0\leq x\leq 1$. Then there exist isometries $U,V,S,T\in\mathbb{M}_{2n,n}$ such that
$|A\oplus B|^{2}=U\,|A\nabla_{x}B|^{2}U^{*}+V\,|B\nabla_{x}A|^{2}V^{*}+x(1-x)\Bigl\{\,S\,|A-B|^{2}S^{*}+T\,|A-B|^{2}T^{*}\Bigr\}.$
Taking traces in Theorem 1.1 yields
$\|A\oplus B\|_{2}^{2}=\|A\nabla_{x}B\|_{2}^{2}+\|B\nabla_{x}A\|_{2}^{2}+2x(1-x)\|A-B\|_{2}^{2}$
for $0\leq x\leq 1$. Since the right-hand side of that trace identity is a polynomial in $x$, it extends to all $x\in\mathbb{R}$. This motivates the following question.
Question 1.2 ([12, Question 2.3]).
Does Theorem 1.1 hold for every $x\in\mathbb{R}$?
We first answer Question 1.2 in the negative: the restriction $0\leq x\leq 1$ in Theorem 1.1 cannot be removed and therefore is sharp (see Theorem 1.3 below).
Theorem 1.3.
Fix $x\in\mathbb{R}\setminus[0,1]$. Then Theorem 1.1 fails for this $x$ and some $A,B\in\mathbb{M}_{n}$. In fact, a counterexample exists already for $n=1$.
1.2. Symmetric moduli and three related questions
For two Hermitian matrices $A,B\in\mathbb{M}_{n}$, we write $A\leq B$ if BAB-A is positive semidefinite. While for $A,B\in\mathbb{M}_{n}$ one has the triangle inequality
$\|A+B\|_{1}\leq\|A\|_{1}+\|B\|_{1},$
simple examples show that the inequality
$|A+B|\leq|A|+|B|$
need not hold in the positive semidefinite order. However, Thompson [20] observed that there exist unitaries $U,V\in\mathbb{M}_{n}$ such that
$|A+B|\leq U|A|U^{*}+V|B|V^{*},$ (1.1)
which, upon taking traces, yields the triangle inequality as a consequence. Such “unitary-orbit” inequalities have since been generalized and extended to a variety of settings; see the surveys by Bourin, Harada and Lee [7, 9].
In this paper, we generalize (1.1) to the setting of rectangular matrices (see Lemma 5.2) in order to address Question 1.11.
Recall the Cartesian decomposition of a matrix $Z\in\mathbb{M}_{n}$:
$Z=\operatorname{Re}Z+i\,\operatorname{Im}Z,\qquad\operatorname{Re}Z:=\frac{Z+Z^{*}}{2},\quad\operatorname{Im}Z:=\frac{Z-Z^{*}}{2i}.$
Note that $|\operatorname{Re}Z|^{2}+|\operatorname{Im}Z|^{2}=\left(|Z|^{2}+|Z^{*}|^{2}\right)/2$. Bourin and Lee [12, Corollary 4.3] obtained the following unitary-orbit estimate for the quadratic symmetric modulus, which can be regarded as a matrix analogue of the scalar inequality $|z|\leq|a|+|b|$ for a complex number $z=a+ib$.
@@ 1.2. Symmetric moduli and three related questions
They also posed the following question concerning the optimality of the exponent $2$.
Question 1.5 ([12, Question 4.6]).
@@
Our second result answers Question 1.5 affirmatively in every noncommutative dimension $n\geq 2$ by constructing a $2\times 2$ counterexample for $p>2$.
@@
For $n=1$, the statement of Question 1.5 is different: for a complex number $z$, the left-hand side of (1.2) equals $|z|$ for every $p>0$, while the right-hand side equals $|\Re z|+|\Im z|\geq|z|$. Hence (1.2) holds for all $p>0$ in dimension 11. The necessity $p\leq 2$ is therefore a genuinely noncommutative phenomenon requiring $n\geq 2$.
@@
Remark 1.8.
In [12, Corollary 4.7], the indices on the right-hand side appear with a typographical slip. Starting from the operator inequality in [12, Corollary 4.3] and applying the Weyl inequality for singular values,
$\mu_{1+j+k}^{\downarrow}(A+B)\leq\mu_{1+j}^{\downarrow}(A)+\mu_{1+k}^{\downarrow}(B)\qquad(A,B\geq 0),$
one obtains the corrected bound
$\mu_{1+j+k}^{\downarrow}\!\left(\sqrt{\frac{|Z|^{2}+|Z^{*}|^{2}}{2}}\right)\leq\mu_{1+j}^{\downarrow}(\operatorname{Re}Z)+\mu_{1+k}^{\downarrow}(\operatorname{Im}Z).$ (1.3)
Accordingly, the natural corrected version of [12, Question 4.8] asks whether there exists $Z\in\mathbb{K}$ such that for every $p>2$ one has
$\mu_{1+j+k}^{\downarrow}\!\left(\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\right)>\mu_{1+j}^{\downarrow}(\operatorname{Re}Z)+\mu_{1+k}^{\downarrow}(\operatorname{Im}Z)$
for some integers $j,k\geq 0$.
Question 1.9 ([12, Question 4.8]).
@@
$\mu_{1+j+k}^{\downarrow}\!\left(\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\right)>\mu_{1+j}^{\downarrow}(\operatorname{Re}Z)+\mu_{1+k}^{\downarrow}(\operatorname{Im}Z)$ (1.4)
@@
We answer Question 1.9 in the affirmative.
@@
There exists $Z\in\mathbb{K}$ and fixed integers $j,k\geq 0$ such that for every $p>2$, (1.4) holds. In our construction one may take $(j,k)=(2,0)$.
Let
$|Z|_{\mathrm{sym}}:=\frac{|Z|+|Z^{*}|}{2},\qquad|Z|_{\mathrm{qsym}}:=\left(\frac{|Z|^{2}+|Z^{*}|^{2}}{2}\right)^{1/2}$
be the arithmetic symmetric modulus or the quadratic symmetric modulus of $Z\in\mathbb{M}_{n}$, respectively.
Now, we turn to a related problem raised by Bourin and Lee concerning Thompson-type triangle inequalities for symmetric moduli, compared with (1.1).
Question 1.11 ([12, Question 4.9]).
Denote by $|Z|_{\mathrm{(q)sym}}$ either the arithmetic symmetric modulus or the quadratic symmetric modulus of $Z\in\mathbb{M}_{n}$. Let $X,Y\in\mathbb{M}_{n}$. Does the Thompson triangle type inequality
$|X+Y|_{\mathrm{(q)sym}}\leq U\,|X|_{\mathrm{(q)sym}}\,U^{*}+V\,|Y|_{\mathrm{(q)sym}}\,V^{*}.$ (1.5)
hold for some unitary $U,V\in\mathbb{M}_{n}$ ?
Our next theorem provides a complete answer to Question 1.11: the inequality (1.5) fails for the arithmetic symmetric modulus but holds for the quadratic symmetric modulus.
Theorem 1.12.
Let $X,Y\in\mathbb{M}_{n}$.
  1. (i)
    In general, there do not exist unitaries $U,V$ such that
    $|X+Y|_{\mathrm{sym}}\leq U|X|_{\mathrm{sym}}U^{*}+V|Y|_{\mathrm{sym}}V^{*}.$
  2. (ii)
    There exist unitaries $U,V$ such that
    $|X+Y|_{\mathrm{qsym}}\leq U|X|_{\mathrm{qsym}}U^{*}+V|Y|_{\mathrm{qsym}}V^{*}.$
1.3. Matrix analogues of Euler’s quadrilateral identity
@@
We recall that an isometry is a rectangular matrix $W\in\mathbb{M}_{m,n}$ such that $W^{*}W=I_{n}$. Bourin and Lee [12] suggested exploring isometry/unitary-orbit refinements of Euler-type identities. Our following three results give isometry-orbit refinements of (1.6).
@@
Naturally, we ask the following question.
Question 1.18.
Does there exist isometries $U_{1},U_{2},U_{3},U_{4}\in\mathbb{M}_{3n,n}$ such that
$\quad+U_{3}\,|B|^{2}\,U_{3}^{*}+U_{4}\,|C|^{2}\,U_{4}^{*}\ ?$
Unfortunately, the answer to Question 1.18 is negative. In fact, take $A=B=C=I_{n}$. Then
Hence the desired identity would become
$4I_{3n}=9U_{1}U_{1}^{*}+U_{2}U_{2}^{*}+U_{3}U_{3}^{*}+U_{4}U_{4}^{*}.$
$4I_{3n}\geq 9U_{1}U_{1}^{*}.$
But $U_{1}U_{1}^{*}$ is an orthogonal projection (of rank $n$), so $\|9U_{1}U_{1}^{*}\|_{\infty}=9$ while $\|4I_{3n}\|_{\infty}=4$, a contradiction. Therefore such isometries cannot exist in general.
1.4. Sharp Clarkson–McCarthy type inequalities on Euler’s identity
For a Hilbert space $\mathcal{H}$, let $S_{p}$ denote the Schatten $p$-class. For $p>0$, we say that $X\in S_{p}$ if $X$ is a compact operator on $\mathcal{H}$ whose singular values $(\mu_{k}^{\downarrow}(X))_{k\geq 1}$ satisfy $\sum_{k\geq 1}\mu_{k}^{\downarrow}(X)^{p}<\infty$ . In this case, the Schatten pp-(quasi)norm is defined by
$\|X\|_{p}:=\Big(\sum_{k\geq 1}\mu_{k}^{\downarrow}(X)^{p}\Big)^{1/p}.$
(When $0<p<1$, $\|\cdot\|_{p}$ is a quasi-norm.)
@@ 1.4. Sharp Clarkson–McCarthy type inequalities on Euler’s identity
Building on the Euler operator identity (1.6), we obtain the following sharp Clarkson–McCarthy type inequalities.
Theorem 1.19.
Let $A,B,C\in S_{p}$. Then for $p\geq 2$,
$\|A+B\|_{p}^{p}+\|B+C\|_{p}^{p}+\|C+A\|_{p}^{p}\leq 2^{p-2}\big(\|A+B+C\|_{p}^{p}+\|A\|_{p}^{p}+\|B\|_{p}^{p}+\|C\|_{p}^{p}\big).$
For $0<p\leq 2$, the inequality is reversed. Moreover, the constant $2^{p-2}$ is optimal: equality holds for all $p>0$ by taking $A=B=-C\neq 0$.
As a complement to Theorem 1.19, numerical evidence suggests the following sharp inequality. For a weaker bound with coefficient $3^{\frac{p}{2}-1}$, see Theorem 7.10.
Conjecture 1.20.
Let $A,B,C\in S_{p}$. Then for $p\geq 2$,
$\|A+B+C\|_{p}^{p}+\|A\|_{p}^{p}+\|B\|_{p}^{p}+\|C\|_{p}^{p}\leq\frac{3^{p-1}+1}{2^{p}}\left(\|A+B\|_{p}^{p}+\|B+C\|_{p}^{p}+\|C+A\|_{p}^{p}\right).$
For $0<p\leq 2$, the inequality is reversed. Moreover, the constant $\frac{3^{p-1}+1}{2^{p}}$ is optimal: equality holds for all $p>0$ by taking $A=B=C\neq 0$.
Classical Hlawka-type inequalities compare norms (or their powers) of vector sums with those of their pairwise sums. A standard form of Hlawka’s inequality [17] asserts that if $V$ is an inner product space and $x,y,z\in V$, then
$\|x+y\|+\|y+z\|+\|z+x\|\ \leq\ \|x\|+\|y\|+\|z\|+\|x+y+z\|,$
where $\|\cdot\|$ denotes the norm induced by the inner product; equivalently, the inequality holds for any norm satisfying the parallelogram identity. Numerous refinements and variants replace $\|\cdot\|$ by powers $\|\cdot\|^{p}$ and/or adjust the constants.
Theorem 1.19 and Conjecture 1.20 can be viewed as a Schatten–$p$ (i.e. noncommutative $L_{p}$) analogue: it features a sharp $p$-dependent constant and exhibits a change in behavior at $p=2$, as is typical for inequalities governed by the uniform convexity/concavity of $L_{p}$ spaces.
As a further refinement of Theorem 1.19, we establish the following sharp mixed $\ell_{q}$$\ell_{p}$ Clarkson–McCarthy type inequalities.
Theorem 1.21.
Let $A,B,C\in S_{p}$. Then for $1<p\leq 2$ ,
$\|A+B+C\|_{p}^{q}+\|A\|_{p}^{q}+\|B\|_{p}^{q}+\|C\|_{p}^{q}$ $\leq 2^{\,1-\frac{q}{p}}\,\Big(\|A+B\|_{p}^{p}+\|B+C\|_{p}^{p}+\|C+A\|_{p}^{p}\Big)^{q/p},$
where $q$ is the conjugate exponent of $p$. For $p\geq 2$, these two inequalities are reversed. Moreover, the constant $2^{1-\frac{q}{p}}$ is optimal: equality holds for all $p>1$ by taking $A=B=-C\neq 0$.
Remark 1.22.
Setting C=BC=-B in Theorems 1.19 and 1.21 recovers the classical Clarkson–McCarthy inequality of [16]. For a complete characterization of the validity of Clarkson–McCarthy type inequalities, see the author’s paper [21, Corollary 1.14] and the following Theorem 1.23 (also see [21, Theorem 1.16]). For further results and developments on Clarkson–McCarthy type inequalities, we refer the reader to [5, 6, 9, 10, 15, 21, 22].
We now record the following mixed-norm inequality (Theorem 1.23), which strengthens the result in the author’s earlier work [21, Theorem 1.16]. Unlike the approach in [21], which applies the three-lines theorem to a suitable trace functional and then invokes a duality argument, we present a new proof based on a vector-valued Riesz–Thorin type complex interpolation scheme.
Theorem 1.23 (Mixed $\ell_{q}(S_{p})$$\ell_{p}(S_{p})$ estimates).
Let $U=[u_{ij}]\in\mathbb{M}_{s,t}$ and set
$\mu:=\max_{1\leq i\leq s,\ 1\leq j\leq t}|u_{ij}|.$
For $1<p<\infty$, let $q$ be the conjugate exponent of $p$, i.e. $\frac{1}{p}+\frac{1}{q}=1$. Define the linear map on $t$-tuples
$T_{U}:\ \ell_{p}^{t}(S_{p})\longrightarrow\ell_{q}^{s}(S_{p}),\qquad T_{U}(Z_{1},\dots,Z_{t}):=\Big(\sum_{j=1}^{t}u_{1j}Z_{j},\ \dots,\ \sum_{j=1}^{t}u_{sj}Z_{j}\Big).$
We will regard $T_{U}$ as an operator between $\ell_{p}$- and $\ell_{q}$-sums when equipped with the corresponding norms.
  1. (i)
    Assume that $U$ is a contraction, i.e. $\|U\|_{\infty}\leq 1$ (in particular, this holds if $U^{*}U=I_{t}$ or $UU^{*}=I_{s}$). Then for $1<p\leq 2$,
    $\Bigg(\sum_{i=1}^{s}\Big\|\sum_{j=1}^{t}u_{ij}Z_{j}\Big\|_{p}^{q}\Bigg)^{1/q}\ \leq\ \mu^{\,\frac{2}{p}-1}\,\Bigg(\sum_{j=1}^{t}\|Z_{j}\|_{p}^{p}\Bigg)^{1/p},$ (1.8)
    and the inequality is reversed for $p\geq 2$ if $U^{*}U=I_{t}$.
  2. (ii)
    Assume in addition that $U^{*}U=I_{t}$ (i.e. $U$ is an isometry). Then for $1<p\leq 2$,
    $\Bigg(\sum_{j=1}^{t}\|Z_{j}\|_{p}^{q}\Bigg)^{1/q}\ \leq\ \mu^{\,\frac{2}{p}-1}\,\Bigg(\sum_{i=1}^{s}\Big\|\sum_{j=1}^{t}u_{ij}Z_{j}\Big\|_{p}^{p}\Bigg)^{1/p},$ (1.9)
    and the inequality is reversed for $p\geq 2$.
Theorem 1.23 has several interesting consequences. For example, Audenaert–Kittaneh’s conjecture [1, Section 8.1, Conjecture 7] (proved by the author in [22]) states that if $A_{1},\ldots,A_{n}\in S_{p}$, then for $1<p\leq 2$,
$\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}\ \leq\ n\left(\sum_{i=1}^{n}\|A_{i}\|_{p}^{p}\right)^{q/p},$ (1.10)
where $q$ is the conjugate exponent of $p$; for $p\geq 2$ the inequality is reversed. Moreover, the constant $n$ is optimal: equality holds for all $1<p<\infty$ by taking $A_{1}=\ldots=A_{n}\neq 0$. This inequality can be recovered from Theorem 1.23 by choosing the isometry
$U=\frac{1}{\sqrt{n}}\begin{pmatrix}1&1&\cdots&1\\ 1&-1&\cdots&0\\ 1&0&\ddots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&-1\end{pmatrix}\in\mathbb{M}_{\frac{n(n-1)+2}{2},\,n},$
whose rows encode the combinations $\sum_{i=1}^{n}A_{i}$ and $A_{i}-A_{j}$. Moreover, Theorem 1.23 yields the following sharp reverse-type companion to (1.10). For more details, see Subsection 7.4.
@@
Let $A_{1},\ldots,A_{n}\in S_{p}$. Then for $1<p\leq 2$,
$\sum_{i=1}^{n}\|A_{i}\|^{q}_{p}\leq n^{-q/p}\left(\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{p}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{p}\right)^{\frac{q}{p}},$ (1.11)
where $q$ stands for the conjugate exponent of $p$. For $p\geq 2$, (1.11) is reversed. Moreover, the constant $n^{-q/p}$ is optimal: equality holds for all $1<p<\infty$ by taking $A_{1}=\ldots=A_{n}\neq 0$.
As further consequences of Euler’s identity (1.6), we next derive sharp Clarkson–McCarthy type inequalities for general unitarily invariant norms, together with a unitary-orbit refinement.
Theorem 1.25.
@@
[…]
For $0<p\leq 2$, the inequality is reversed. Moreover, the constant $2^{p-2}$ is optimal: equality holds for all $p>0$ by taking $A=B=-C\neq 0$.
Theorem 1.26.
Let $A,B,C\in\mathbb{M}_{n}$. Then for $p\geq 2$, there exist unitaries $U_{1},U_{2},U_{3}\in\mathbb{M}_{n}$ such that
$U_{1}|A+B|^{p}U_{1}^{*}+U_{2}|B+C|^{p}U_{2}^{*}+U_{3}|C+A|^{p}U_{3}^{*}\leq 2^{p-2}\big(|A+B+C|^{p}+|A|^{p}+|B|^{p}+|C|^{p}\big).$
For $0<p\leq 2$, the inequality is reversed. Moreover, the constant $2^{p-2}$ is optimal: equality holds for all $p>0$ by taking $A=B=-C\neq 0$.
Organization of this paper. Section 2 gives a negative answer to the extension problem for the matrix-weighted parallelogram law: for each $x\notin[0,1]$ we construct a counterexample, proving Theorem 1.3. In Section 3, we prove Theorem 1.6 and show that the range $p\leq 2$ in the Bourin–Lee unitary-orbit inequality (Theorem 1.4) is optimal by exhibiting an explicit $2\times 2$ obstruction for every $p>2$. Section 4 establishes Theorem 1.10 via a compact-operator construction that forces the failure of the corresponding singular-value inequality for all $p>2$, and hence confirms the optimality of (1.3). In Section 5, we resolve the Thompson-type triangle problem for symmetric moduli (Theorem 1.12). Finally, Sections 68 develop isometry-orbit refinements of an Euler operator identity (Theorems 1.15, 1.16, 1.17), derive Clarkson–McCarthy type inequalities, and draw further consequences for unitarily invariant norms and singular values.
2. Proof of Theorem 1.3
Proof of Theorem 1.3.
Let $n=1$ and take scalar matrices
$A=1,\qquad B=\frac{x-1}{x}\quad(\text{note that }x\neq 0\text{ since }x\notin[0,1]).$
Then
$A\nabla_{x}B=(1-x)A+xB=(1-x)+x\cdot\frac{x-1}{x}=0,$
while
$B\nabla_{x}A=(1-x)B+xA=(1-x)\frac{x-1}{x}+x=\frac{2x-1}{x}\neq 0,$
because $x\notin[0,1]$ implies $x\neq\tfrac{1}{2}$. Moreover,
$A-B=1-\frac{x-1}{x}=\frac{1}{x}\neq 0.$
Assume, for contradiction, that Theorem 1.1 holds for this $x$. Since $x\notin[0,1]$, we have x(1x)<0x(1-x)<0. With $A\nabla_{x}B=0$, the identity in Theorem 1.1 would reduce to
$|A\oplus B|^{2}=V\,|B\nabla_{x}A|^{2}V^{*}+x(1-x)\Bigl\{\,S\,|A-B|^{2}S^{*}+T\,|A-B|^{2}T^{*}\Bigr\},$
for some isometries $V,S,T\in\mathbb{M}_{2,1}$. Moving the last term to the left and using x(1x)>0-x(1-x)>0, we obtain
$|A\oplus B|^{2}+(-x(1-x))\Bigl\{\,S\,|A-B|^{2}S^{*}+T\,|A-B|^{2}T^{*}\Bigr\}=V\,|B\nabla_{x}A|^{2}V^{*}.$
Here $|A\oplus B|^{2}=\operatorname{diag}(|A|^{2},|B|^{2})=\operatorname{diag}\!\bigl(1,|(x-1)/x|^{2}\bigr)$ is positive definite, and the bracketed term is positive semidefinite; hence the whole left-hand side is positive definite, thus has rank $2$.
On the other hand, since $V\in M_{2,1}$ is an isometry, $VV^{*}$ is a rank-one projection; moreover $|B\nabla_{x}A|^{2}>0$. Therefore the right-hand side $V\,|B\nabla_{x}A|^{2}V^{*}$ has rank $1$.
This is impossible. Hence Theorem 1.1 cannot hold for this $x$. Since $x\in\mathbb{R}\setminus[0,1]$ was arbitrary, the claim follows. ∎